OpenStudy (anonymous):

four times the lesser of two consecutive even integers is 12 less than twice the greater number. find the integers

OpenStudy (anonymous):

you just take it apart and write it down as you go: "Four times.." so ..... 4 x ( ? ) " ....the lesser of 2 consecutive even integers" you think .. ok 2 consecutive even integers have to be something and then something + 2, so choose a letter for the something, let's say s, so now we have "four times the lesser of two consecutive integers .... 4 x s ... or 4s for short, so ... 4s " .... is 12 less than" .. something , which means to equal something you have to add 12 to it. so : 4s +12 = ? "...twice the greater number" .... the greater number is s+2 like we said before so : 4s +12 = 2(s + 2) Now simplify : 4s + 12 = 2s + 4 ... subtract 2s from both sides, then subtract 12 from both sides: 2s = -8 ... divide by 2: s = -4

OpenStudy (anonymous):

Thank you so much !! You made it so thorough so I could understand :D

OpenStudy (anonymous):

Np :)